OpenStudy (anonymous):

a^2+5a+4/a^3 divided by a^2+3a+2/a^2-2a don't know where to start :/ algebra 1 @sleepyjess

OpenStudy (loser66):

\[\frac{a^2+5a+4}{a^3}\div \frac{a^2+3a+2}{a^2-2a}\]

OpenStudy (anonymous):

yep, that makes it more clear. Do you have time to explain this to me? I don't really know the steps

OpenStudy (loser66):

Do you know how to factor? Do you know \(\dfrac{a}{b}\div \dfrac{c}{d}= \dfrac{a}{b}*\dfrac{d}{c}\)??

OpenStudy (anonymous):

i do somewhat, no where near an expert on it though..

OpenStudy (loser66):

If you don't know , now you do. Apply the division property above (the second one), show me how you turn your problem from division to multiplication, please

OpenStudy (anonymous):

\[a^2+5a+4/a^3 \div a^2-2a/a^2+3a+2\]

OpenStudy (anonymous):

i don't know how to post it like you did

OpenStudy (anonymous):

but hopefully you see what I'm trying to do

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

multiplication, not division, right? the terms are ok

OpenStudy (anonymous):

oh right, yea multiplication

OpenStudy (loser66):

\(\dfrac{a^2+5a+4}{a^3}*\dfrac{a^2-2a}{a^2+3a+2}\) So far so good

OpenStudy (loser66):

Now, factor. The first term: numerator: a^2 + 5a+4 = ??? factor it, please

OpenStudy (anonymous):

correct me if I'm wrong but is this where i could use the FOIL method?

OpenStudy (loser66):

Whatever, no matter what method is, you can use FOIL, Discriminant, or else. I need its factor.

OpenStudy (anonymous):

urg, i may need to go review factoring and then get back to you :/ unless your up to give a quick explanation on how to do it. I just can't remember when using foil, which numbers are paired with each other

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

ok, as a gift, I give you the full solution, try to figure out how to do it by yourself next time. \(\dfrac{(a+1)(a+4)}{aa^2}*\dfrac{a(a-2)}{(a+1)(a+2)}\\=\dfrac{\cancel{(a+1)}(a+4)}{\cancel{a}a^2}*\dfrac{\cancel{a}(a-2)}{\cancel{(a+1)}(a+2)}\)

OpenStudy (anonymous):

thanks for that gift :P so thats the answer ? or the outcome of factoring?

OpenStudy (loser66):